OpenStudy (anonymous):

Find the derivative of y=lnsqrt(x^2-3)

OpenStudy (anonymous):

start with \[y=\frac{1}{2}\ln(x^2-3)\] and then it will be much easier

OpenStudy (anonymous):

using \[\frac{d}{dx}[\ln(f(x))]=\frac{f'(x)}{f(x)}\]

OpenStudy (anonymous):

you can just about do it in your head. try it

OpenStudy (anonymous):

Thanks! Yup I got there. If I let u = x^2 -3, does dy/du = 1/2u?

OpenStudy (anonymous):

oh, we haven't learnt that formula yet.

OpenStudy (anonymous):

what "formula"?

OpenStudy (anonymous):

this ddx[ln(f(x))]=f′(x)f(x) formula

OpenStudy (anonymous):

ignore the 1/2 out front, get \[\frac{2x}{x^2-3}\] instantly, then divide by 2 and get \[\frac{x}{x^2-3}\]

OpenStudy (apoorvk):

Use the chain rule. \[\frac{d(ln\sqrt{x^2 - 3})}{dx} = \frac{1}{2}\frac{d(ln\sqrt{x^2-3})}{dx}\] now, derivative of ln(x) is 1/x. but here, the argument of the log is (x^2 - 3). So, use chain rule.

OpenStudy (anonymous):

it is not a "formula" it the chain rule

OpenStudy (anonymous):

the derivative of log of x is one over x the derivative of log of something is one over something times the derivative of something that is the chain rule in action more succinctly written as \[\frac{d}{dx}[\ln(f(x))]=\frac{f'(x)}{f(x)}\]

OpenStudy (anonymous):

@apoorvk think there is a typo there. square root should be gone after the equal sign

OpenStudy (apoorvk):

Oops sorry! thanks @satellite73 , yes the sqrt would be gone after the equal sign. should read: \[\frac{d(ln\sqrt{x^2 - 3})}{dx} = \frac{1}{2}\frac{d~ln{(x^2-3})}{dx}\]

OpenStudy (anonymous):

Hmm.. ok I understand the chain rule but when I used to u to sub x^2 -3, y = 0.5lnu, d(0.5linu)/du = 1/2u ... which isn't the top number that you can get from doing f'(x)/f(x)

OpenStudy (anonymous):

OH I get it now! Thank you both!!